• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 16:05:06 on May16,2007

    Entry number 202311

    keyword=day shift summary

    Day shift summary

    Shift workers: J. P. Chen (SL), R. Gilman (TO)

    8:00-9:13 production data taking 105 uA on He3
    9:13-9:27 move target to dummy.
    9:27-10:27 40 uA on 20 cm dummy.
    10:27-10:38 move target to He3.
    10:38 back to 105 uA on He3.
    12:45-13:00 MCC took beam away.
    13:00-16:00 production data taking ~105uA on He3

    run 3497 8:05-8:25, 1.25M, 105 uA on He3.
    run 3498 8:25-8:52, 1.25M, 105 uA on He3.
    run 3499 8:52-9:13, 1.25M, 105 uA on He3.

    run 3500 9:27-9:50, 0.32M, 40 uA on 20cm dummy. Did not change PS factors.
    run 3501 9:50-10:27. 1.25M, 40 uA on 20 cm dummy.

    run 3502 10:38-10:59,1.25M, 105 uA on He3.
    run 3503 10:59-11:20,1.25M, 105 uA on He3.
    run 3504 11:21-11:40,1.25M, 105 uA on He3.
    run 3505 11:40-11:59,1.25M, 105 uA on He3.
    run 3506 11:59-12:17,1.25M, 105 uA on He3.
    run 3507 12:17-12:38,1.25M, 105 uA on He3.
    run 3508 12:38-12:50,1.25M, 105 uA on He3.
    run 3509 13:01-13:21,1.25M, 105 uA on He3.
    run 3510 13:21-13:40,1.25M, 105 uA on He3.
    run 3511 13:41-14:01,1.25M, 104 uA on He3.
    run 3512 14:01-14:20,1.25M, 104 uA on He3.
    run 3513 14:21-14:41,1.25M, 104 uA on He3.
    run 3514 14:41-14:59,1.25M, 104 uA on He3.
    run 3515 15:00-15:19,1.25M, 104 uA on He3.
    run 3516 15:19-15:37,1.25M, 104 uA on He3.
    run 3516 15:37-15:55,1.25M, 104 uA on He3.
    run 3516 15:55-? 104 uA on He3. still on-going at 16:00.